OpenStudy (malia_halvorson):

I just need to know the steps on how to solve this: If the area of a trapezoid is 40 cm squared, the height is 4 cm and the bottom base is 12 cm, what is the length of the top base?

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

Area of trapezoid \[\Large A = h*\frac{b_1+b_2}{2}\]

jimthompson5910 (jim_thompson5910):

b1 and b2 are the parallel bases h is the height

OpenStudy (malia_halvorson):

My online teacher never taught me to divide the bases by 2

jimthompson5910 (jim_thompson5910):

how about multiplying the sum by 0.5?

jimthompson5910 (jim_thompson5910):

maybe your teacher gave you something like this? \[\Large A = h*0.5*(b_1+b_2)\]

OpenStudy (malia_halvorson):

My classes are online so I never interact with him. He just gives us questions and videos to watch but the videos don't really explain it. But i would just plug in what I already know right?

jimthompson5910 (jim_thompson5910):

yes `If the area of a trapezoid is 40 cm squared, the height is 4 cm and the bottom base is 12 cm` A = 40 h = 4 b1 = x b2 = 12

OpenStudy (malia_halvorson):

40=4* b1+12/2 ??

jimthompson5910 (jim_thompson5910):

\[\Large A = h*\frac{b_1+b_2}{2}\] \[\Large 40 = 4*\frac{x+12}{2}\] \[\Large 40 = 4*\frac{1}{2}*(x+12)\] \[\Large 40 = 2*(x+12)\] \[\Large 40 = 2x+24\] solve for x

OpenStudy (malia_halvorson):

x=8

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (malia_halvorson):

That was easier than I thought. I understood it more thank you

jimthompson5910 (jim_thompson5910):

no problem